How many grams of F are in 415 g of CaF2?
1
Expert's answer
2015-10-28T11:57:06-0400
M (Ca) = 40.1
M (2 F) = 38.0
M = 78.1
%F = 38/78.1
%F = 48.7
m(F) = 0.487 * 415 = 202.1

Answer: 202.1